OpenStudy (anonymous):

The perimeter of a rectangular ranch is 400 m. Find the dimensions of the ranch that will contain the greatest area.

OpenStudy (anonymous):

Wow every time i go to answer a question that i am incapable of answering you answer for me :)

OpenStudy (anonymous):

so could be 50+50+100+100

OpenStudy (anonymous):

could also be all 75

OpenStudy (anonymous):

I am horrible at this stuff even though it looks very simple

OpenStudy (anonymous):

500 or 150

OpenStudy (anonymous):

do the 500 one

OpenStudy (anonymous):

so 50+50+100+100

OpenStudy (anonymous):

I know it's just a simple question but I just don't know how to start it.

OpenStudy (anonymous):

it's an optimization problem

OpenStudy (anonymous):

use the derivative to find the local maximum which in this case, the largest area

OpenStudy (anonymous):

uhmmm, how? i don't know that @11calcBC

OpenStudy (anonymous):

set x = length and y = width, u know that 2x + 2y = 400, so x =200 - y

OpenStudy (anonymous):

implicit differentiation...

OpenStudy (anonymous):

2(200 - y) + 2y = 400 400 - 2y + 2y = 400 400 = 400

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

well...it's right but plugging it in isn't solving it lol

OpenStudy (anonymous):

u plug 200 - y into the area, so u get A(x) = (200-y)y

OpenStudy (anonymous):

which you simplify to 200y - y^2

OpenStudy (anonymous):

take the derivative of that, which is 200 - 2y

OpenStudy (anonymous):

set that equal to 0, get ur critical point, which is 100 for y when you solve for it. remember to find the x value, which is 2x + 2(100) = 400, which is also equal to 100. also note that a number is the greatest when it's two values are the same

OpenStudy (anonymous):

ur maximum area will then be 100 x 100 or 10,000

OpenStudy (anonymous):

thank you so much for letting me understand and for helping me out with this :) @11calcBC :)

OpenStudy (anonymous):

my pleasure :D